### v1

# Parafermions in Moiré Minibands

Abstract
Moiré materials provide a remarkably tunable platform for topological and strongly correlated quantum phases of matter. Very recently, the first zero field Abelian fractional Chern insulators (FCIs) have been experimentally demonstrated and it has been theoretically predicted that non-Abelian states with Majorana fermion excitations may be realized in the nearly dispersionless minibands of these systems. Here we provide telltale evidence in terms of low-energy quantum numbers, spectral flow and entanglement spectra for the even more exotic possibility of moiré-based non-Abelian FCIs exhibiting Fibonacci parafermion excitations.
Introduction.—In the last years, moiré materials have been established as an accessible and highly tunable platform for exploring strongly correlated topological phases of matter [1, 2, 3, 4, 5, 6, 7]. Most notably, specific van der Waals heterostructures consisting of twisted layers of graphene or transition metal dichalcogenides have been theoretically predicted [8, 9, 10, 11, 12, 13, 14, 15] and experimentally observed [16, 17, 18, 19, 20, 21, 22] to host fractional Chern insulators (FCIs)—lattice analogs of fractional quantum Hall (FQH) states that can exist in the absence of a magnetic field [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. So far, the FCI states that have been unambiguously identified in experiments are topologically equivalent to Abelian hierachical FQH states [34, 35, 36, 37] hosting Abelian anyon excitations [38, 39]. While the possibility of exploring anyon physics at zero magnetic field and moderate temperatures is already remarkable, the pursuit of non-Abelian quasiparticles—which should appear as elementary excitations in more exotic FCI states—constitutes a more ambitious and potentially rewarding venture with prospects for fault-tolerant topological quantum computation [40].
Recent experimental signatures [41], suggested to arise from non-Abelian topological order at moiré filling fractions with even denominator, have sparked a series of theoretical works addressing the stability of Moore-Read (MR) states in moiré systems [42, 43, 44, 45, 46, 47, 48, 49]. Strengthening the theoretical evidence based on the ground state degeneracy, which can correspond to either the MR phase or charge density waves (CDWs), further analysis counting quasi-hole excitations in entanglement spectra has shown that MR states could indeed be stable in moiré materials [47, 42, 48]. Unfortunately, the braiding operations allowed by the emergent Majorana excitations in MR states cannot generate any arbitrary unitary transformation and would thus be insufficient for universal quantum computation [40]. This issue can, however, be overcome by exploiting parafermions with richer braiding statistics, which have been predicted to appear in FQH systems [50], FQH–superconductor heterostructures [51], and certain non-Hermitian systems [52], although their experimental realization remains elusive. A natural step forward in the context of moiré FCIs is thus to search for Fibonacci parafermions appearing in the $\mathbb{Z}_{3}$ Read-Rezayi (RR) state [50]—which is characterized by the clustering of composite fermions into triplets, as opposed to the pairing in the MR phase.
In this Letter, we provide conclusive evidence showing that the non-Abelian RR FCI at 3/5 filling can be realized in a moiré system. In particular, we perform many-body exact diagonalization on double twisted bilayer graphene (dTBG) and obtain the expected number of degenerate ground states appearing at momenta fulfilling the generalized Haldane statistics. Furthermore, we verify the persistence of the many-body energy gap upon insertion of an external magnetic flux, and, more strikingly, we show that the state counting in the low-energy sector of the particle-cut entanglement spectrum (PES) exactly matches the number of allowed quasi-hole excitations. In addition, we find that the RR phase is destroyed when the average quantum metric of the considered flat band approaches asymptotically the ideal value for the second Landau level (LL), highlighting both the usefulness and the limitations of this quantity as a heuristic indicator for the emergence of non-Abelian phases. Finally, we show that the RR phase is less stable at $2/5$ filling but still present—as reflected in the quasi-particle excitations accessed through the hole entanglement spectrum—, signalling the robustness of moiré-based parafermion FCIs.
Figure 1: Evidence for moiré parafermions at $\nu=3/5$. (a) Low-lying many-body energy spectrum, (b) spectral flow, and (c) $N_{A}=5$ particle-cut entanglement spectrum for a $N_{\text{s}}=20$-site system at band filling $\nu=3/5$ and $\gamma=3.75$. The corresponding results for $N_{\text{s}}=25$ sites are shown in (d)-(f). The 10-fold nearly-degenerate ground states are marked by red dots. In the PES, the number of states below the red solid line is $14404$ for $N_{\text{s}}=20$ and $51255$ for $N_{\text{s}}=25$, both matching the quasi-hole excitation counting for Read-Rezayi states detailed in the SM. (g) Structure factor and (h) pair-correlation function of the Read-Rezayi ground states for $N_{\text{s}}=25$. The black lines in (h) denote the unit cell of the system. The spanning vectors for the considered systems are $\mathbf{T}_{1}=(3,4)$ and $\mathbf{T}_{2}=(-2,4)$ for $N_{\text{s}}=20$, and $\mathbf{T}_{1}=(5,0)$ and $\mathbf{T}_{2}=(0,5)$ for $N_{\text{s}}=25$.
Setup.—We consider dTBG as an exemplary moiré system with the valley- and spin-polarized Hamiltonian [53]
$H_{0}(\bf r)=\begin{pmatrix}u_{1}I&D_{1}^{\dagger}(\bf r)&0&0\\ D_{1}(\bf r)&u_{2}I&\gamma I&0\\ 0&\gamma I&u_{3}I&D_{2}^{\dagger}(\bf r)\\ 0&0&D_{2}(\bf r)&u_{4}I~{}\end{pmatrix},$ (1)
[…]
In order to tackle the many-body problem of interacting electrons, we diagonalize $H_{0}(\bf r)$ in the basis of Bloch states and project the electron-electron interactions onto the considered nearly-flat band, obtaining $H_{\text{int}}=\frac{1}{2}\sum_{\bf k_{1}k_{2}k_{3}k_{4}}V_{\bf k_{1}k_{2}k_{3}k_{4}}c_{\bf k_{1}}^{\dagger}c_{\bf k_{2}}^{\dagger}c_{\bf k_{3}}c_{\bf k_{4}}.$ Here, the matrix element $V_{\bf k_{1}k_{2}k_{3}k_{4}}$ ensures momentum conservation and contains single-particle form factors [60, 8] as well as the 2D bare Coulomb potential $V({\bf q})=~{}\frac{1}{2A\epsilon|{\bf q}|}$ with the average dielectric constant of the system $\epsilon=10$ [61]. Next, we obtain the low-lying energy spectrum of many-body states by exact diagonalization of the spin- and valley-polarized interaction Hamiltonian $H_{\text{int}}$. In particular, we consider a periodic system of $N_{\text{e}}$ electrons in $N_{\text{s}}$ moiré unit cells with the filling $\nu=N_{\text{e}}/N_{\text{s}}=3/5,2/5$.
Read-Rezayi parafermion states.—We start by considering a fractionally filled dTBG with $\nu=3/5$ and a relatively large coupling $\gamma=3.75$ between the two TBG sheets. Performing exact diagonalization on systems with size $N_{\text{s}}=20$ and $N_{\text{s}}=25$, we obtain 10 nearly-degenerate ground states separated from the excited states by an energy gap, cf. red dots in Fig. 1(a),(d). The splitting of these states is generically expected on the relatively small finite size systems accessible to exact diagnalization, but should go away in the large system limit. Moreover, we find that the ground states do not flow into excited states upon insertion of a magnetic flux, confirming the presence of a many-body gap, cf. Fig. 1(b),(e). The 10-fold ground state degeneracy is in agreement with that expected from RR states in the thin-torus limit, where the ground states at filling $\nu=k/(kM+2)$ are formed by $k$ particles occupying $kM+2$ consecutive sites and every two particles being separated by at least $M$ sites [62, 50, 63]. For fermionic Read-Rezayi states with $k=3$ and $M=1$, i.e. $\nu=3/5$, two distinct configurations are then expected, namely $1110011100\cdots$ and $1101011010\cdots$, which together with their translation-invariant partners result in a 10-fold degenerate ground state. Furthermore, the calculated ground states in Fig. 1(a),(d) appear at center-of-mass momenta matching those expected from the aforementioned thin-torus configurations.
While the approximate ground state degeneracy and the spectral flow is consistent with the RR phase, it could also correspond to weakly entangled states such as CDWs—as has been shown for candidate MR states [47]. Thus, more information is needed to conclusively determine whether the ground states in Fig. 1(a),(d) correspond to RR states. To this end, we compute the PES of the ground states [64, 23]. The PES is obtained by dividing the many-body system into $A$ and $B$ subsystems consisting of $N_{A}$ and $N_{B}=N_{\text{e}}-N_{A}$ particles, and then calculating the eigenvalues of $-\mathrm{log}\rho_{A}$, where $\rho_{A}=\mathrm{tr}[~{}\frac{1}{N_{\text{d}}}~{}\sum_{i=1}^{N_{\text{d}}}~{}\ket{\Psi_{i}}\bra{\Psi_{i}}]$ is the reduced density matrix of $A$. Here, $\rho_{A}$ carries crucial information about the quasi-hole excitations in the $N_{\text{d}}$-fold degenerate ground states |$\ket{\Psi_{i}}$, which are characteristically distinct for Abelian FCIs, non-Abelian FCIs, and CDWs. In particular, a gap in the PES is expected, below which the number of states exactly matches the amount of quasi-hole excitations allowed by the specific quantum phase of the system.
In Fig. 1(c),(f) we show the PES with $N_{A}=5$ for the two considered system sizes. In addition, we mark with a red line the entanglement energy below which the number of states matches exactly the amount of allowed quasi-hole excitations (the analytical counting is detailed in the SM [55]). For the smaller system ($N_{\text{s}}=20$), the dense low-energy sector is followed by a region of sparsely distributed states resembling a gap (around the red line in Fig. 1(c)), below which the state counting is similar to that expected for the RR phase. Remarkably, the PES for the larger system ($N_{\text{s}}=25$) shows a clear entanglement gap with the low-energy state counting being exactly equal to the number of quasi-hole excitations in the RR phase. This observation serves as smoking-gun evidence that the considered system is in the RR phase.
To further clarify the nature of the phase, we calculate the structure factor and pair-correlation function averaged over the tenfold quasi-degenerate ground states, which provide insights on the crystalline or liquid character of the system. Here, the structure factor helps measure if the ground states exhibit a preferred charge order and is defined as $S(\mathbf{q})=\frac{1}{N_{\text{e}}}(\langle\rho(\mathbf{q})\rho(-\mathbf{q})\rangle-\delta_{\mathbf{q},0}N_{\text{e}}\langle\rho(\mathbf{q})\rangle),$ with $\rho(\mathbf{q})$ being the density operator projected onto the nearly flat band. On the other hand, the real space pair-correlation function reads $g(\mathbf{r})=~{}\frac{1}{N^{2}_{\text{e}}}(\braket{n(\mathbf{r})n(\mathbf{0})}-\delta_{\mathbf{r},0}\braket{n^{2}(\mathbf{r})})$, where $n(\mathbf{r})$ is the real-space density operator, and can be obtained via the Fourier transform of $S(\mathbf{q})$. For liquid phases, $g(\mathbf{r})$ approaches a constant for large $\mathbf{r}$, while it shows a periodic arrangement of charges for CDWs. As shown in Fig. 1(g)-(h), both the structure factor and the pair-correlation function indicate a liquid behavior—concretely, no prominent peaks in $S(\mathbf{q})$ and no CDW unit cells appear in $g(\mathbf{r})$.
Figure 2: Stability of the parafermion phase and quantum geometry. The many-body energy spectrum gap for non-Abelian $\mathbb{Z}_{3}$ RR states at $\nu=3/5$ (red dotted lines, right axis) as a function of the coupling strength $\gamma$ between twisted layers is shown for system sizes $N_{\text{s}}=20$ (squares) and $N_{\text{s}}=25$ (circles) together with the averaged quantum geometry (blue line, left axis).
Stability of parafermion FCIs and quantum geometry.—Having demonstrated the existence of non-Abelian RR states, we now explore the stability of such topological phases in the parameter space associated with the quantum geometry, which serves as a heuristic indicator of the analogy between the partially filled moiré band and the LLs in two-dimensional electron gases [49]. For the $n$-th LL, the quantum geometry and the Berry curvature satisfy [57] $\text{tr}[g(\mathbf{k})]=(2n+1)|\Omega(\mathbf{k})|$, with $n=0,1,2...$. It is then generally believed that a moiré band is expected to mimic the respective LL physics if this relation is approximately satisfied. Following this principle, it has been recently shown that both the Laughlin-like zero-energy modes [10] and non-Abelian MR state can be realized in a moiré flat band with a (nearly) ideal quantum geometry [42, 43, 44, 45, 46, 48]. As mentioned before, the average quantum geometry $\chi$ of the considered flat band in dTBG increases gradually from $1$ to $3$ as $\gamma$ is increased from 00 to infinity (see blue line in Fig. 2). The parameter $\gamma$ then provides an ideal tuning knob to study the transition between the physics of the lowest LL and the second LL.
In the strong coupling limit, $\gamma\rightarrow\infty$, the targeted band should in principle emulate second LL physics, as it acquires an ideal quantum geometry $\chi=3$ and the band dispersion becomes exactly flat. However, while the RR states are energetically competitive in the first excited Landau level [58, 59], the many-body low-lying energy spectrum gap shown in Fig. 2 indicates their absence as $\gamma\rightarrow\infty$. Instead, the $\mathbb{Z}_{3}$ RR appears at an intermediate coupling strength, ranging from $\gamma\approx 3$ to $4.5$. Here, the energy gap is defined as $\Delta_{E}=E_{\text{e}}-E_{\text{g}}$ with $E_{\text{e}}$ being the $11$th lowest energy and $E_{\text{g}}$ being the $10$th lowest energy. These findings strongly suggest that the quantum geometry criterion is not sufficient to diagnose the non-Abelian topological phases. It nevertheless gives some intuition of where exotic states may occur—the RR states do in fact form in a region where the quantum geometry resembles that of the first excited rather than that of the lowest LL. We note that there are alternative ways of making moiré minibands that are close but not identical to the first excited LL such as considering other magic angles [65], adding a “negative” magnetic field [66] or periodic strain [67], and considering bands further away from charge neutrality [42, 46, 48, 43].
In the weak coupling limit $\gamma\rightarrow 0$, where the system approaches to two decoupled chiral TBG sheets, the average quantum geometry $\chi\rightarrow 1$ shown in Fig. 2 suggests that the targeted flat band resembles the lowest LL, where Abelian hierarchy states at band filling $\nu=3/5$ and its particle-hole partner $\nu=2/5$ [35, 36, 37] are indeed present (see the SM [55] for more details).
Particle-hole asymmetry and hole entanglement spectrum.—We now search for the possibility of RR states at electronic band filling $\nu=2/5$, which in ideal LLs correspond to the particle-hole conjugate of the $\nu=3/5$ RR phase. It has been shown that, in general, particle-hole symmetry is broken in FCIs and moiré systems [68, 8], a fact that can bee seen as a consequence of a fluctuating quantum geometry  [69, 70]. In line with this asymmetry, we do not find a clear many-body energy gap of the RR states for filling $\nu=2/5$ at the system sizes considered, suggesting that the RR phase is less stable for $\nu=2/5$ than for $\nu=3/5$ (see Fig. 3(a)). In addition, although the tenfold quasi-degenerate states are still present at the correct momentum sector, their corresponding PES does not show a gap with the correct quasi-hole counting (see the SM [55]). This, however, should not be surprising: since the $\nu=2/5$ RR phase can be understood as the RR phase for holes at 3/5 hole-filling, adding holes to the system directly increases the ground-state energy as the 4-body interaction can no longer be minimized.
Instead, the $\nu=2/5$ RR phase allows a certain number of quasi-particle excitations without increasing the ground-state energy. This is reflected in the hole entanglement spectrum (HES), where the many-body system is divided into $A$ and $B$ subsystems with $N_{A}$ and $N_{B}=N_{\text{s}}-N_{\text{e}}-N_{A}$ now denoting the number of holes. Importantly, the HES (Fig. 3(b)) displays an entanglement gap below which the number of states matches the quasi-particle counting of the 3/5 hole-filling (i.e. $\nu=2/5$) RR phase, indicating that the calculated states indeed correspond to the non-Abelian RR phase. From a topological perspective, this then implies that the nature of the RR states remains largely unaffected by the renormalization of the hole dispersion that is expected to arise from fluctuations in the quantum geometry [68], even though these fluctuations may alter the ground state energies as suggested by the absence of a gap in Fig. 3(a).
Figure 3: Moiré parafermions and competing states at $\nu=2/5$. (a) Low-lying energy spectrum at band filling $\nu=2/5$ and (b) the respective hole entanglement spectrum (HES) for $\gamma=3.75$ in a system with size $N_{\text{s}}=25$. In the HES, the number of states below the first entanglement gap (red line) is 51255, matching the quasi-particle counting of RR states at hole-filling 3/5.
Discussion.—In this work we have demonstrated the emergence of Fibonacci parafermions in the exemplary moiré system dTBG. While a quantitatively realistic material prediction would require a consideration of finite bandwidth [71] and band mixing effects [72, 73, 74] as well as reliable estimates of system parameters (which in many cases remain controversial), our results are very encouraging for several reasons.
First, that an unscreened Coulomb interaction is sufficient to obtain these states is quite remarkable given that, for the simplest non-Abelian FCIs, fine-tuned interactions are necessary even in toy models  [75, 76].
Second, although there is suggestive evidence for the possibility of RR parafermion quantum Hall states in Landau levels at $\nu=12/5$ and $13/5$ [58, 59], competing orders make the case less compelling than that of MR states states at $\nu=5/2$. This situation may be opposite in the moiré context: at half filling of moiré bands charge density waves [15] and the anomalous Hall crystal [77] are additional fierce competitors with no direct LL analogues, while presumably no similarly competitive non-LL states exist at $\nu=3/5$ due to a lack of natural ordering patterns.
Third, mimicking Landau levels is not the optimal strategy for finding RR parafermion states. Fig. 2 instead suggests that the most optimal bands have an average quantum geometry intermediate between the lowest and first excited Landau levels. Effects of a fluctuating geometry are unavoidable and generally favour competing states [70, 69], yet our findings suggest that the tunability of moiré materials and hence the average geometric properties are more important for the realization of RR parafermion FCIs.
Fourth, although we find the RR parafermion states to be energetically favoured in a moderate parameter regime (cf. Fig. 2), it is striking that, even in cases when it is not apparent from the small system sizes avaliable whether the RR states will be competitive for a macroscopic number of particles, the entanglement spectrum shows telltale signatures of the RR states (cf. Fig. 3) suggesting that these states may prevail in a larger parameter regime.
Finally, we note that Fibonacci parafermions are arguably more desirable and elusive than Majorana fermions for fundamental reasons [78]: unlike the Majoranas, they are not possible to realize in effectively non-interacting Hermitian models [52] and they potentially serve as building blocks for universal quantum computing [78]. These aspects have motivated the study of highly complex and technically extremely challenging setups such as FQH–superconductor heterostructures [51]. In contrast, moiré heterostructures offer a simple, versatile and tunable platform which does in principle not even require an external magnetic field. We hope that these observations will serve as inspiration for both the theoretical and experimental pursuit of parafermions emerging in moiré minibands.
Acknowledgements.
Acknowledgements.— We acknowledge useful discussions and related collaborations with Ahmed Abouelkomsan, Liang Fu, Zhao Liu, Aidan Reddy and Donna Sheng. This work was supported by the Swedish Research Council (VR, grant 2018-00313), the Wallenberg Academy Fellows program of the Knut and Alice Wallenberg Foundation (2018.0460) and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine.